# Resolvent-Based Optimisation for Invariant Flows

### v1

@@ 2.2 Galerkin Projection
Ashtari & Schneider (2023) proposed a methodology for equilibrium that avoids this issue by solving for both the update to the velocity field, $\bm{u}$, and the pressure field, $p$, simultaneously in a coupled fashion using the IM method, which ensures that both $\divergence{\bm{r}}=0$ and $\left.\bm{r}\right|_{y=\pm 1}=0$ are satisfied. The same treatment is provided for $\bm{r}$ and $q$ in a staggered approach, providing a physically consistent evolution of both $\bm{u}$ and $\bm{r}$. However, as mentioned, this method was developed to solve for equilibria ($\partialderivative*{t}=0$), and it is not obvious how this would work for temporally varying fields because the method used to obtain compatible velocity and pressure fields is designed as an update method for some time-stepping scheme.
With the difficulties in evolving the variational dynamics while satisfying all the constraints now clear, we propose to use a Galerkin projection onto an orthogonal basis that satisfies the boundary conditions and the incompressibility constraints. Some setup is required to arrive at the final procedure that yields a valid computation of the residuals and gradient that obey the constraints. First, the velocity field is expanded into a sum over a basis as follows
$\bm{u}\left(x,\,y,\,z,\,t\right)=\bm{u}_{b}+\sum_{i=1}^{\infty}\sum_{\bm{k}\in\mathbb{Z}^{3}}a_{\bm{k}i}\bm{\psi}_{\bm{k}i}\left(y\right)e^{i\bm{k}\cdot\bm{\xi}},$ (11)
where $\bm{k}=\left(k_{x}\quad k_{z}\quad k_{t}\right)^{\top}$ is the integer wavenumber vector, and $\bm{\xi}=\left(\alpha x\quad\beta z\quad\omega t\right)^{\top}$ is the scaled direction vector. The coefficients $\alpha=2\pi/L_{x}$, $\beta=2\pi/L_{z}$, and $\omega=2\pi/T$ are determined by the domain size and period of the flow, and each represent the smallest frequency in each of their respective direction that can be accommodated in the finite space. The modes $\bm{\psi}_{\bm{k}i}=\bm{\psi}_{\bm{k}i}\left(y\right)$ are divergence-free and obey the no-slip boundary conditions
@@
where $\delta_{ij}$ is the Kronecker delta, and $(\cdot)^{\dagger}$ denotes the conjugate transpose. The operator $\gradient_{\bm{k}}\dotproduct=i\alpha k_{x}+\partial/\partial y+i\beta k_{z}$ is the divergence operator in spectral/Fourier space. To account for the inhomogeneous boundary conditions on $\bm{u}$, the steady base flow $\bm{u}_{b}=\bm{u}_{b}\left(y\right)$ is introduced into (11). The sum over the set of modes $\bm{\psi}_{\bm{k}i}$ is taken over a countably infinite set of modes for every frequency $\bm{k}$ to form a complete basis for the optimisation space $\mathcal{P}_{T}$. In practice, when the problem is discretised or when forming a reduced-order model of the dynamics, this sum is truncated to a finite number denoted with $M\in\mathbb{N}$. The key observation is that for any combination of coefficients $a_{\bm{k}i}\in\mathbb{C}$ the velocity field is incompressible and obeys the no-slip boundary conditions, restricting the velocity field to exist only within a finite-dimensional sub-space of the complete function space.
@@
$a_{\bm{k}i}=\frac{1}{2TL_{z}L_{x}}\left\langle\bm{u}-\bm{u}_{b},\,e^{-i\bm{k}\dotproduct\bm{\xi}}\bm{\psi}_{\bm{k}i}\right\rangle_{\Omega_{t}}$ (15)
@@
Figure 3 is a flow diagram for the data over a single iteration of a given optimisation loop, to compute both $\mathcal{R}$ and $\partialderivative*{\mathcal{R}}{a_{\bm{k}i}}$. The computation begins with the coefficients $a_{\bm{k}i}$ that represent a given periodic and incompressible velocity field that obeys the periodicity and no-slip boundary conditions, relative to a given set of modes $\bm{\psi}_{\bm{k}i}$ and base flow $\bm{u}_{b}$. The residual coefficients $s_{\bm{k}i}$ are then obtained by computing all the terms in (5) excluding the pressure gradient $\gradient{p}$, the result of which is then projected onto the basis as in (17). The global residual can be obtained directly from $s_{\bm{k}i}$. To obtain the residual gradient, the velocity and residual coefficients are expanded back into physical space to facilitate the computation of all the terms on the right-hand side of (8) excluding the adjoint pressure gradient $\gradient{q}$. Finally the result is projected using (18) to obtain $\partialderivative*{\mathcal{R}}{a_{\bm{k}i}}$. With both $\mathcal{R}$ and its gradient, the chosen optimisation algorithm can be used to update the coefficients $a_{\bm{k}i}$ to reduce $\mathcal{R}$, with convergence being determined by the size of $\mathcal{R}$. The process summarised here allows the direct computation of the required residual and gradient without resorting to solving a PPE, avoiding the issues related to the boundary conditions. The frequency gradient $\partialderivative*{\mathcal{R}}{\omega}$ does not require any extra special treatment as once the appropriate local residual $\bm{r}$ has been determined then (10) can be computed directly.
It should be noted that the spatial derivative terms in (18) and (19) are not explicitly computed in terms of the coefficients $a_{\bm{k}i}$ and $s_{\bm{k}i}$ here. These expressions would depend on the explicit form of the modes $\bm{\psi}_{\bm{k}i}$ as could be seen by substituting the expansions (11) and (16) into (18), and thus a general statement cannot be provided. In addition, the computation of the nonlinear terms in projected space using $a_{\bm{k}i}$ and $s_{\bm{k}i}$ would be in the form of several convolution sums, as can be seen in Barthel et al. (2021). These sums scale very poorly with the degrees-of-freedom of the system, and for this reason a pseudo-spectral approach is taken. The velocity and residual are expanded back into their full physical space forms where the nonlinear terms are computed after which they are then projected back onto the modes. This avoids the heavy computation of convolutional sums in the reduced-order space, replacing them with pointwise multiplication in physical space. Truncating the number of modes used in the expansion/projection steps of the procedure to a potential small number of dynamically significant modes forms a low-order model for the dynamics. This low-order model is characterised by the shape of the modes $\bm{\psi}_{\bm{k}i}$ and the number of modes retained. The primary effects of this truncation are discussed in section 6.
@@ 3 Resolvent Analysis
To implement the projected variational optimisation methodology, a method to generate the modes $\bm{\psi}_{\bm{k}i}$ with the desired properties (12)-(14) is required. This choice is non-trivial with many possible solutions. For example, SPOD could be used to generate a set of modes from large data sets, or global stability analysis can derive modes directly from the Navier-Stokes equations. Here resolvent analysis is used to generate the desired modes, and for completeness we provide an overview of resolvent analysis following closely the formulation provided in Sharma (2019).
@@
$\bm{u}\left(x,\,y,\,z,\,t\right)=\bm{u}_{b}\left(x,\,y,\,z\right)+\bm{u}^{\prime}\left(x,\,y,\,z,\,t\right),$ (21)
where the base flow here is distinct from the one used in (11). The standard choice in resolvent analysis is to take $\bm{u}_{b}$ to be the turbulent mean. The choice taken in this work is different for primarily pragmatic reasons and is discussed further in sections 5 and 6.
At this point it is useful to define the Leray projector $\mathbb{P}\bm{u}=\bm{u}-\gradient\bm{\Delta}^{-1}(\divergence{\bm{u}})$ Temam (1984). This operator takes an instantaneous velocity field and orthogonally projects onto the divergence-free subspace, and can be derived from Helmholtz decomposition of the velocity field. Applying the Leray projector to (1a) and then substituting in the decomposition (21) provides an evolution equation for the fluctuations
@@
The operator $\mathcal{N}\left(\bm{u}\right)$ is the Navier-Stokes operator, given by the right-hand side of (1a). The operator $\mathcal{L}_{\bm{u}_{b}}$ is the linearised Navier-Stokes operator evaluated at the base flow $\bm{u}_{b}$. Finally, $\bm{f}\left(\bm{u}^{\prime}\right)=-\left(\bm{u}^{\prime}\dotproduct\gradient\right)\bm{u}^{\prime}$ is the nonlinear term for the fluctuations $\bm{u}^{\prime}$. Physically $\bm{f}$ the nonlinear feedback that transports energy between scales and sustains any unsteady motion in the flow. The addition of the Leray projector allows us to ignore the pressure gradient terms in the operators $\mathcal{N}$ and $\mathcal{L}_{\bm{u}_{b}}$, as well as not needing to explicitly represent the continuity equation in (22).
@@
ranked in order of the associated singular values $\sigma_{i}\geq\sigma_{i+1}\geq 0$, for all $i\in\mathbb{N}$. The modes $\bm{\psi}_{\bm{k}i}$ and $\bm{\phi}_{\bm{k}i}$ are the left and right singular modes and form a complete basis for the range (response) and domain (input) of the resolvent for every $\bm{k}\in\mathbb{Z}^{3}\setminus\left\{0\right\}$, respectively.
To actually find exact solutions as discussed in section 2, a basis for the mean $\bm{k}=0$ is also required. The resolvent in (26) is only technically defined using the fluctuation equation at all non-zero frequencies, $\bm{k}\neq 0$, however the resolvent operator can regardless be evaluated at $\bm{k}=0$. The exact physical interpretation of the basis used for the Galerkin projection is less important as its ability to provide a valid orthonormal basis in the space of divergence-free and no-slip fields. The literature discussing the evaluation of the resolvent at the mean frequencies and wavenumbers is extremely sparse, since it is not typically required as the mean is used as an input. The exception is Mons & Marquet (2021) which uses resolvent of the RANS for the purpose of data assimilation.
Since the range of the resolvent is the space of divergence-free fluctuation velocity fields that obey the desired boundary condition, the left singular modes, $\bm{\psi}_{\bm{k}i}$, also called the response modes, have these desired properties. Thus, the response modes $\bm{\psi}_{\bm{k}i}$ can be used as the basis for the Galerkin projection introduced in section 2.2.
The rate at which the singular values $\sigma_{i}$ decay determines how accurately the decomposition (27) can be represented with only a finite number of the modes. It has been observed for fluid flows that the singular values decay rather rapidly (McKeon & Sharma, 2010), meaning a low-rank approximation to the resolvent operator can be constructed with a partial sum of (27). This reduces the dimensionality of the amplification mechanisms in (26), only keeping the most significant contributions.
@@ 4 General Features of Rotating Plane Couette Flow
We now introduce RPCF, defining the governing equations for this flow, as well as its general features and normal solutions at various regimes. Figure 4 is a diagram of the flow geometry between two parallel plates moving in opposite directions in a reference rotating around the spanwise direction.
The governing equations for this flow are given by
@@
in addition to the periodicity boundary conditions in the streamwise and spanwise directions. In this work we make an additional assumption that the flow is streamwise-independent ($\partial/\partial x=0$) reducing it to a 2-Dimensional 3-Component (2D3C) formulation. The primary reason to use a 2D3C formulation of the flow is to reduce the dimensionality of the problem making the solutions easier to find while retaining many of the key features of the original dynamics. This constraint in fact makes the flow exactly analogous to Rayleigh-Benard convection as shown in (Eckhardt et al., 2020). The analogy between rotating shear flows and convective thermal flows has been known for some decades (Chandrasekhar, 1961; Yih, 1965), and has been used to describe Taylor-Couette flows in turbulent regimes in relation to the Rayleigh Benard convection (Eckhardt et al., 2007). The nature of some steady and streamwise-independent solutions to the RPCF at low Reynolds numbers (before bifurcating into more complex structures) is discussed in Nagata (2013); Nagata et al. (2021).
Figure 4: Schematic for the rotating plan Couette flow with nomenclature.
The streamwise-independence means that the domain size of the flow is given solely by the half-channel height $h$ and the spanwise length $L_{z}$, with the aspect ratio defined as the ratio $\gamma=L_{z}/h$. In this work an aspect ratio of $\gamma=4$ is used exclusively, which is equivalent to setting $L_{z}=8$. The flow is characterised by two non-dimensional parameters, the Reynolds number $\mathit{Re}$ and Rotation number $\mathit{Ro}$ defined as
@@
where $U_{w}$ is the wall speed, $\nu$ is the kinematic viscosity, and $\Omega$ is dimensional rotation rate. The Rotation number is the ratio between the characteristic rotation and inertial forces of the system. In this work only the case of the rotation being oriented to point out of the page is considered. This is known as anti-cyclonic rotation and acts to produce linear instabilities at finite Reynolds numbers (Tsukahara et al., 2010), as opposed to standard plane Couette flow that is linearly stable for all finite Reynolds numbers (Daviaud et al., 1992). In general, the most interesting parameter range for the Rotation number is $0<\mathit{Ro}<1$ as this is range that destabilises the flow. For any value outside this range for $\mathit{Ro}$, the flow is actually made more linearly stable by the presence of system rotation (Lezius & Johnston, 1976; Hiwatashi et al., 2007).
@@
The flow starts as with a stable laminar solution that is identical that of plane Couette flow. Linear stability analysis shows the boundary of stability for the laminar flow has the following relationship
$\mathit{Re}=\sqrt{\frac{107}{\mathit{Ro}\left(1-\mathit{Ro}\right)}},$ (33)
@@
given in Lezius & Johnston (1976). This initial bifurcation of the fully 3D flow leads to a streamwise flow, and so it is also a valid relation for the 2D3C case. This implies that the laminar flow is most unstable for $\mathit{Ro}=0.5$, which is the rotation number of choice for this work. At $\mathit{Re}\approx 20.7$ the laminar state bifurcates and becomes unstable. A new stable equilibrium solution is born out of this bifurcation which has the characteristic streamwise-independent rolls that are present throughout the total set of regimes accessible by varying $\mathit{Re}$.
@@
Figure 5: Bifurcation diagram of RPCF over a range of Reynolds number and $\mathit{Ro}=0.5$ showing the transition from the stable laminar solution to turbulent flow.
We characterise the transition behaviour by analysing the kinetic energy extrema, defined as $\mathcal{K}(t)=\frac{1}{2}\norm{\bm{u}(t)}_{\Omega}$, of the DNS data over a range of Reynolds numbers at $\mathit{Ro}=0.5$. The results are shown in figure 5. This and any further DNS results shown in this work were obtained using a custom solver originally developed for Lasagna et al. (2016) using a vorticity-streamfunction formulation of the flow. The data for figure 5 was obtained by simulating a random initial condition at the given Reynolds number for long enough to ensure the simulation has converged to the final attractor, typically taking of the order of thousands of time units. A full analysis of the transitionary behaviour of 2D3C RPCF is beyond the scope of this work. However, there are some key features in figure 5 that it will be useful to further discuss. As the Reynolds number is further increased above the laminar bifurcation the streamwise rolls grow in strength. This continues until the equilibrium flow becomes unstable enough to yield to a periodic flow. The exact bifurcation that occurs is rather complex, with a stable equilibrium and periodic solution seemingly existing at the same time for certain ranges of Reynolds numbers. Once the Reynolds number has reached about $\mathit{Re}=350$, the DNS converge towards a single periodic solution. The spatial structure of this stable periodic solution does not differ much from the equilibrium, with the majority of the kinetic energy being stored in large streamwise rolls. This stable periodic orbit gradually gives way to more complex non-periodic behaviour. Eventually, the flow reaches a point where the flow has become sufficiently complex to be considered turbulent, with a broadband spectrum and notable separation of scales.
@@
The initial bifurcation from the laminar state occurs at the expected value predicted from (33) and the kinetic energy of the new stable equilibrium grows quickly. Notably, instead of the single equilibrium being traced continuously as the Reynolds number increases the favoured attractor the DNS settles on changes. These new branches represent alternative stable equilibrium solutions. The equilibrium solutions at $\mathit{Re}=50$ for each branch labelled in figure 5 are shown in figure 6. These solutions are labelled as S1, S2, and S3, as depicted in figure 5. The solutions of each branch differ primarily in the number of streamwise rolls. The solutions have kinetic energies of $\mathcal{K}\approx 0.65$, $\mathcal{K}\approx 0.58$, and $\mathcal{K}\approx 0.39$, for S1-3 respectively. The denser the rolls the lower the kinetic energy as a result of the higher dissipation rates at the same Reynolds number. Also of note is that the higher the Reynolds number the more favoured the denser rolling structures become, likely as a result of the previous branch not being able to sustain the required dissipation rates to produce a stable equilibrium. It should be noted that all of these solutions are linearly stable at $\mathit{Re}=50$ as the DNS converged towards each of them as long as the initial condition is sufficiently close. The primary reason for some the branches, especially S3, not being observed at this Reynolds number is likely due to a small basin of attraction at $\mathit{Re}=50$.
Figure 6: Equilibrium solutions for branch shown in figure 5 in the Reynolds range of $20\lessapprox\mathit{Re}\lessapprox 100$.

### v2

@@ 2.2 Galerkin Projection
Ashtari & Schneider (2023) proposed a methodology specific to equilibrium solutions that avoids this issue by solving for both the update to the velocity field, $\bm{u}$, and the pressure field, $p$, simultaneously in a coupled fashion using the Influence Matrix (IM) method, which ensures that both $\divergence{\bm{r}}=0$ and $\left.\bm{r}\right|_{y=\pm 1}=0$ are satisfied. The same treatment is provided for $\bm{r}$ and $q$ in a staggered approach, providing a physically consistent evolution of both $\bm{u}$ and $\bm{r}$. However, this method was developed to solve for equilibria ($\partialderivative*{t}=0$), and it is not obvious how this would work for temporally varying fields because the method used to obtain compatible velocity and pressure fields is designed as an update method for some time-stepping scheme.
With the difficulties in evolving the gradient-based optimisation while satisfying all the constraints now clear, we propose to use a Galerkin projection onto an orthogonal basis that satisfies the boundary conditions and the incompressibility constraints. Some setup is required to arrive at the final procedure that yields a valid computation of the residuals and gradient that obey the constraints. First, the velocity field is expanded into a sum over a basis as follows
$\bm{u}\left(x,\,y,\,z,\,t\right)=\bm{u}_{\text{BC}}(y)+\sum_{i=1}^{\infty}\sum_{\bm{k}\in\mathbb{Z}^{3}}a_{\bm{k}i}\bm{\psi}_{\bm{k}i}\left(y\right)e^{i\bm{k}\cdot\bm{\xi}},$ (11)
where $\bm{k}=\left(k_{x}\quad k_{z}\quad k_{t}\right)^{\top}$ is the integer wavenumber vector, and $\bm{\xi}=\left(\alpha x\quad\beta z\quad\omega t\right)^{\top}$ is the scaled direction vector. The coefficients $\alpha=2\pi/L_{x}$, $\beta=2\pi/L_{z}$, and $\omega=2\pi/T$ are determined by the domain size and period, and each represent the smallest frequency in each of their respective direction that can be accommodated in the finite space. The modes $\bm{\psi}_{\bm{k}i}=\bm{\psi}_{\bm{k}i}\left(y\right)$ are divergence-free and obey the no-slip boundary conditions
@@
where $\delta_{ij}$ is the Kronecker delta, and $(\cdot)^{\dagger}$ denotes the conjugate transpose. The operator $\gradient_{\bm{k}}\dotproduct=i\alpha k_{x}+\partial/\partial y+i\beta k_{z}$ is the divergence operator in spectral/Fourier space. To account for the inhomogeneous boundary conditions on $\bm{u}$ in (2), the steady base flow $\bm{u}_{\text{BC}}\left(y\right)$ is introduced into (11). The sum over the set of modes $\bm{\psi}_{\bm{k}i}$ is taken over a countably infinite set of modes for every frequency $\bm{k}$ to form a complete basis for the optimisation space $\mathcal{P}_{T}$. This includes the mean mode $\bm{k}=0$ as this is part of the solution that needs to determined during the optimisation. In practice the sums in (11) are truncated to be finite. The key observation is that for any combination of coefficients $a_{\bm{k}i}\in\mathbb{C}$ the velocity field is incompressible and obeys the no-slip boundary conditions, restricting the velocity field to exist only within a finite-dimensional sub-space of the complete function space.
@@
$a_{\bm{k}i}=\frac{1}{2TL_{z}L_{x}}\left\langle\bm{u}-\bm{u}_{\text{BC}},\,e^{-i\bm{k}\dotproduct\bm{\xi}}\bm{\psi}_{\bm{k}i}\right\rangle_{\Omega_{t}}$ (15)
@@
Figure 3 is a flow diagram for the computations over a single iteration of a given optimisation loop, to compute both $\mathcal{R}$ and $\partialderivative*{\mathcal{R}}{a_{\bm{k}i}}$. The computation begins with the coefficients $a_{\bm{k}i}$ that represent a given periodic and incompressible velocity field that obeys the periodicity and no-slip boundary conditions, relative to a given set of modes $\bm{\psi}_{\bm{k}i}$ and base flow $\bm{u}_{\text{BC}}$. The residual coefficients $s_{\bm{k}i}$ are then obtained by computing all the terms in (5) excluding the pressure gradient $\gradient{p}$, the result of which is then projected onto the basis as in (17). The global residual can be obtained directly from $s_{\bm{k}i}$. To obtain the residual gradient, the velocity and residual coefficients are expanded back into physical space to facilitate the computation of all the terms on the right-hand side of (8) excluding the adjoint pressure gradient $\gradient{q}$. Finally the result is projected using (18) to obtain $\partialderivative*{\mathcal{R}}{a_{\bm{k}i}}$. With both $\mathcal{R}$ and its gradient, the chosen optimisation algorithm can be used to update the coefficients $a_{\bm{k}i}$ to reduce $\mathcal{R}$, with convergence being determined by the size of $\mathcal{R}$. The process summarised here allows the direct computation of the required residual and gradient without resorting to solving a PPE, avoiding the issues related to the boundary conditions. The frequency gradient $\partialderivative*{\mathcal{R}}{\omega}$ does not require any extra special treatment as once the appropriate local residual $\bm{r}$ has been determined then (10) can be computed directly.
It should be noted that the spatial derivative terms in (18) and (19) are not explicitly computed in terms of the coefficients $a_{\bm{k}i}$ and $s_{\bm{k}i}$ here. These expressions would depend on the explicit form of the modes $\bm{\psi}_{\bm{k}i}$ as could be seen by substituting the expansions (11) and (16) into (18), and thus a general statement cannot be provided. In addition, the computation of the nonlinear terms in projected space using $a_{\bm{k}i}$ and $s_{\bm{k}i}$ would be in the form of several convolution sums, as can be seen in Barthel et al. (2021). These sums scale very poorly with the degrees-of-freedom of the system, and for this reason a pseudo-spectral approach is taken. The velocity and residual are expanded back into their full physical space forms where the nonlinear terms are computed after which they are then projected back onto the modes. This avoids the heavy computation of convolutional sums in the mode space, replacing them with pointwise multiplication in physical space.
@@ 3 Resolvent Analysis
To implement the projected variational optimisation methodology, a method to generate the modes $\bm{\psi}_{\bm{k}i}$ with the desired properties (12)-(14) is required. This choice is non-trivial with many possible solutions. For example, SPOD could be used to generate a set of modes from large data sets, or global stability analysis can derive modes directly from the Navier-Stokes equations. Here resolvent analysis is used to generate the desired modes which form a complete basis as in (11) and (16). For completeness we provide an overview of resolvent analysis following closely the formulation provided in Sharma (2019).
@@
$\bm{u}\left(x,\,y,\,z,\,t\right)=\bm{u}_{b}\left(y\right)+\bm{u}^{\prime}\left(x,\,y,\,z,\,t\right),$ (21)
where the base flow here is the flow about which the equations will be linearised, serving a distinct role from the base flow $\bm{u}_{\text{BC}}(y)$ used in (11). The standard choice in resolvent analysis is to take $\bm{u}_{b}$ to be the turbulent mean. The choice taken in this work is different for primarily pragmatic reasons and is discussed further in sections 5 and 6.
At this point it is useful to define the Leray projector $\mathbb{P}\bm{u}=\bm{u}-\gradient\bm{\Delta}^{-1}(\divergence{\bm{u}})$ (Temam, 1984). This operator takes an instantaneous velocity field and orthogonally projects onto the divergence-free subspace, and can be derived from Helmholtz decomposition of the velocity field. Applying the Leray projector to (1a) and then substituting in the decomposition (21) provides an evolution equation for the fluctuations
@@
The operator $\mathcal{N}\left(\bm{u}\right)$ is the Navier-Stokes operator, given by the right-hand side of (1a). The operator $\mathcal{L}_{\bm{u}_{b}}$ is the linearised Navier-Stokes operator evaluated at the base flow $\bm{u}_{b}$. Finally, $\bm{f}\left(\bm{u}^{\prime}\right)=-\left(\bm{u}^{\prime}\dotproduct\gradient\right)\bm{u}^{\prime}$ is the nonlinear term for the fluctuations $\bm{u}^{\prime}$. Physically $\bm{f}$ represents the nonlinear feedback that transports energy between scales and sustains any unsteady motion in the flow. The addition of the Leray projector allows us to ignore the pressure gradient terms in the operators $\mathcal{N}$ and $\mathcal{L}_{\bm{u}_{b}}$, as well as not needing to explicitly represent the continuity equation in (22).
@@
ranked in order of the associated singular values $\sigma_{i}\geq\sigma_{i+1}\geq 0$, for all $i\in\mathbb{N}$. The modes $\bm{\psi}_{\bm{k}i}$ and $\bm{\phi}_{\bm{k}i}$ are the left and right singular modes and form a complete basis for the range (response) and domain (input) of the resolvent for every $\bm{k}\in\mathbb{Z}^{3}\setminus\left\{0\right\}$, respectively. Since the range of the resolvent is the space of divergence-free fluctuation velocity fields that obey the desired boundary condition, the left singular modes, $\bm{\psi}_{\bm{k}i}$, also called the response modes, have these desired properties. Thus, the response modes $\bm{\psi}_{\bm{k}i}$ can be used as the basis for the Galerkin projection introduced in section 2.2.
To actually find exact solutions as discussed in section 2, a basis for the mean $\bm{k}=0$ is also required. The resolvent in (26) is only technically defined using the fluctuation equation at all non-zero frequencies, $\bm{k}\neq 0$. However, it was found that the resolvent operator can regardless be evaluated at $\bm{k}=0$ which provided the required modes for the mean, completing the basis required to find exact solution using the expansions (11) and (16). The exact physical interpretation of the basis used for the Galerkin projection is less important as its ability to provide a valid orthonormal basis in the space of divergence-free and no-slip fields.
The rate at which the singular values $\sigma_{i}$ decay determines how accurately the decomposition (27) can be represented with only a finite number of the modes. It has been observed for fluid flows that the singular values decay rapidly (McKeon & Sharma, 2010), meaning a low-rank approximation to the resolvent operator can be constructed with a partial sum of (27). This reduces the dimensionality of the amplification mechanisms in (26), only keeping the most significant contributions.
@@ 4 General Features of Rotating Plane Couette Flow
We now introduce RPCF as the main flow configuration being used as a test case for the variational optimisation methodology. RPCF is defined on the same domain as the general planar wall-bounded flow in section 2.1 with the same Cartesian coordinate system. The channel has a height of $2h$, the top and bottom walls move in opposite directions at a speed of $U_{w}$, and the domain is rotated about the spanwise direction at a rotational rate of $\Omega$. The (non-dimensional) governing equations for this flow are given by
@@
in addition to the periodicity boundary conditions in the streamwise and spanwise directions. The flow is characterised by two non-dimensional parameters, the Reynolds number $\mathit{Re}$ and rotation number $\mathit{Ro}$ defined as
@@
where $\nu$ is the kinematic viscosity. The rotation number is the ratio between the characteristic rotation and inertial forces of the system. In this work, only the rotation direction aligned with the shear is considered. This is known as anti-cyclonic rotation and acts to produce linear instabilities at finite Reynolds numbers (Tsukahara et al., 2010), as opposed to standard plane Couette flow that is linearly stable for all finite Reynolds numbers (Daviaud et al., 1992). In general, the most interesting parameter range for the rotation number is $0<\mathit{Ro}<1$ as linear instabilities in the flow are present in this range. For any value outside this range for $\mathit{Ro}$, the flow is actually made more linearly stable due to the presence of system rotation (Lezius & Johnston, 1976; Hiwatashi et al., 2007).
In this work we make an additional assumption that the flow is streamwise-independent ($\partial/\partial x=0$) reducing it to a 2-Dimensional 3-Component (2D3C) formulation. The primary reason to use a 2D3C formulation is to reduce the dimensionality of the problem making the solutions easier to find while retaining some of the key features of the original dynamics. This constraint in fact makes the flow exactly analogous to 2D Rayleigh-Benard convection as shown in Eckhardt et al. (2020). The analogy between rotating shear flows and convective thermal flows has been known for some decades (Chandrasekhar, 1961; Yih, 1965), and has been used to describe Taylor-Couette flows in turbulent regimes in relation to the Rayleigh Benard convection (Eckhardt et al., 2007). The nature of some steady and streamwise-independent solutions to the RPCF at low Reynolds numbers (before bifurcating into more complex structures) is discussed in Nagata (2013) and Nagata et al. (2021). The streamwise-independence means that the domain size is given solely by the half-channel height $h$ and the spanwise length $L_{z}$, with the aspect ratio defined as the ratio $\gamma=L_{z}/h$. In this work an aspect ratio of $\gamma=4$ is used exclusively.
@@
The flow starts with the stable laminar solution $\bm{u}=y\hat{\imath}$. Linear stability analysis shows that the boundary of stability for the laminar flow has the following relationship
$\mathit{Re}_{\text{crit}}=\sqrt{\frac{107}{\mathit{Ro}\left(1-\mathit{Ro}\right)}},$ (33)
@@
given in Lezius & Johnston (1976). This initial bifurcation leads to a streamwise independent flow, and so it is also a valid relation for the 2D3C case. This implies that the laminar flow is most unstable for $\mathit{Ro}=0.5$, which is the rotation number of choice for this work. At $\mathit{Re}\approx 20.7$ the laminar state bifurcates and becomes unstable. A new stable equilibrium solution is born out of this bifurcation which has the characteristic streamwise-independent rolls that are present throughout the total set of regimes accessible by varying $\mathit{Re}$.
@@
Figure 4: Bifurcation diagram of RPCF over a range of Reynolds number and $\mathit{Ro}=0.5$ showing the transition from the stable laminar solution to turbulent flow. Also plotted are the corresponding kinetic energy extrema of the periodic solution obtained from optimisation at $\mathit{Re}=400$ in section 5.3, denoted with red triangles.
We characterise the route to turbulence by analysing time series extrema of the kinetic energy from DNS data over a range of Reynolds numbers at $\mathit{Ro}=0.5$, where the kinetic energy is defined as $\mathcal{K}(t)=\frac{1}{2}\norm{\bm{u}(t)}_{\Omega}$. All DNS results in this work were obtained using a custom solver (Lasagna et al., 2016) which employs a vorticity–streamfunction formulation, details of which are provided in appendix D. The results are shown in figure 4. The purpose of this analysis is to provide the necessary context for interpreting the results from the variational optimiser. A comprehensive study of the transitionary behaviour of 2D3C RPCF lies beyond the present scope. For each Reynolds number, the DNS was initialised with a new random initial condition and then integrated until it converged to a final attractor, typically requiring thousands of time units. This procedure means that slightly different solutions may be obtained for different runs, depending on the initial condition.
@@
The initial bifurcation from the laminar state occurs at the expected value $\mathit{Re}\approx 20.7$, as predicted from (33). The kinetic energy of the resulting stable equilibrium grows rapidly as $\mathit{Re}$ increases. Notably, instead of a single equilibrium branch being traced continuously, the DNS often settles on different attractors. These correspond to distinct stable equilibrium solutions, examples of which are labelled in figure 4. At $\mathit{Re}=50$, three such branches, S1–S3, are shown in figure 5. They differ primarily in the number of streamwise rolls and exhibit kinetic energies of approximately $\mathcal{K}=0.65$, $0.58$, and $0.39$ for S1, S2, and S3 respectively. The denser the rolls, the lower the kinetic energy, owing to higher dissipation at fixed Reynolds number. As $\mathit{Re}$ increases, denser rolling structures are favoured, likely because sparser structures cannot sustain the required dissipation rates. All of these equilibria are linearly stable at $\mathit{Re}=50$, since the DNS converges to them provided the initial condition lies within their basin of attraction. However, branches such as S3 are less commonly observed at this Reynolds number, likely because their basins of attraction are very small.
As $\mathit{Re}$ increases further, each of these equilibria undergoes bifurcations to periodic solutions, sustained only over limited parameter ranges before seemingly losing stability and no longer acting as an attractor for the DNS. Around $\mathit{Re}=200$, the S1 branch again emerges as the dominant equilibrium, before bifurcating near $\mathit{Re}=350$ to a periodic orbit sustained up to a little above $\mathit{Re}=500$. This orbit retains the large-scale streamwise roll structure of its parent equilibrium. Beyond $\mathit{Re}=500$, the periodicity becomes increasingly irregular, with large amplitude fluctuations and a more broadband spectrum, indicating transition to turbulence. Finally, at the largest Reynolds numbers displayed in figure 4 the flow bifurcates again into a new apparently periodic solution before again transitioning to a chaotic evolution. These remaining bifurcations have not been investigated in any substantial way here.
In addition to the main branches, figure 4 also shows intermittent deviations that appear only over very narrow ranges of Reynolds number. These most likely arise because each DNS run was initialised with a different random condition, leading to variations in the transient dynamics. The precise origin of these deviations has not been investigated here, but they could correspond either to additional stable branches with small basins of attraction, which are only rarely observed, or to trajectories becoming temporarily trapped in transient regions of state space before converging to the dominant attractor. A more systematic exploration of state space would be required to determine their exact nature.
Figure 5: Equilibrium solutions labelled in figure 4 at $\mathit{Re}=50$.